• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 00:02:00 on October 7,2002

    Entry number 87521

    keyword=Shift Summary

    continuing kinematics ep4-3
    16:44 Q1 right tripped.
    Asked MCC for an access
    16:45 Q1 left tripped.
    called the tech on call (Ed) because the two He level are really low (-25% for left, 5% for right)
    17:10: Q1s fixed
    17:27 beam back, 20 uA
    19:12 Vertical scan for the Compton (during run 1180). Beam did NOT move on the target during the scan.
    21:00 Done with kinematics ep4-3. We have a tot of 12.1 M events. Last run at this kinematics: 1182
    We called MCC to ask for the beam to be secured.
    HRS left is now at 1.687 GeV/c
    HRS right is now at 3.064 GeV/c (Q2 and Q3 cycled)
    We moved left HRS to 43.604 degree
    We moved right HRS to 22.318 degree
    We opened the 1.5" and 3" FPP carbon doors
    21:52 We are ready for data taking. Asked for the beam to be back.
    21:54. Beam back (20 uA).
    run 1183: coda crash
    runs 1184-1189: short runs to adjust the prescalers so we are at deadtime < 5% following the instructions. But T5 rate is too big (~500 Hz) for being at deadtime <5% so we run at 8-10%. T1, T2, T3 and T5 are prescaled so that rates are ~10-20 Hz
    first good run for this kinematics: 1190
    22:50 Q1 left and right tripped (exactly the same as at 16:44 and 16:45 He levels are at -25% and 5%). Stop run 1190
    We asked MCC for an access and to reset the Q1s. The Q1 right arm did not reset properly: The power supply display did not clear the ~"earth link failure" (or something like that). So I turned the power off again and did again an hard reset, a soft reset and then the error cleared.
    23:24 Back to beam permit. Q1 right still ramping up.
    23:32 Q1 right back. Asked for the beam but MCC has to finish cycling a magnet.
    23:34. Beam back. Start data taking.